Gravel is being dumped from a conveyor belt at a rate of 20 cubic feet per minute. It forms a pile in the shape of a right circular cone whose base diameter and height are always the same. How fast is the height of the pile increasing when the pile is 19 feet high?
Recall that the volume of a right circular cone with height h and radius of the base r is given by v= 1/3pi(r)^2h


okay so this is what i did and got but it isnt right and i dont know why...

20 = pi/4 3(19)^2 * dH/dt
and i solved for dH/dt..and i got 80/1083pi which is 0.0235 but its wrong...why? thanks for the help!
Answer #1
You said the base diameter and height were the same
then if the height is 1
the radius is 9.5 not 19
answered by: Tiki
Answer #2
V= PI/3 (h)(h/2)^2=PI/12 h^3

dV/dt= PI/12 *3h^2 dh/dt

dh/dt= dV/dt*4/(PIh^2)

dh/dt= 20*4/(PI*19^2)=not your answer.

Take a close look at the difference
answered by: jal
Answer #3
Whoops, pay attention to what bobpursley did.
answered by: Beau
Answer #4
diameter = 19, radius = 9.5
I mean
answered by: Roy Jr